# StyleDrive: Towards Driving-Style Aware Benchmarking of End-To-End Autonomous Driving

### v2

@@ 1 Introduction
As autonomous driving (AD) technologies continue to mature, one of the key factors for translating these systems into real-world products is personalization. Tailoring vehicle behavior to individual user preferences is essential for enhancing user experience, building trust, and fostering long-term adoption Hasenjäger and Wersing (2017). Traditional modular systems have extensively explored personalized strategies, enabling adaptations to specific driving preferences. However, these methods often rely on isolated, scenario-specific adaptations Tian et al. (2022) or unrealistic human-in-the-loop (HITL) simulation Ke et al. (2024), which severely limit their generalization to dynamic, real-world environments. Furthermore, the fragmented nature of modular systems Zhu et al. (2018); Cui et al. (2024); Kou et al. (2025) hinders scalability to massive real-world data. Due to these limitations, personalization remains largely underexplored in end-to-end autonomous driving (E2EAD), where perception, planning, and control are integrated in an unified architecture Codevilla et al. (2018). This gap presents a significant barrier to realizing human-centric AD at scale.
@@
This gap is particularly pressing. As E2EAD systems become increasingly capable and are deployed across diverse, open-world driving environments, aligning vehicle behavior with user preferences becomes essential for enhancing comfort, perceived safety, and long-term user acceptance Speidel et al. (2019); Aledhari et al. (2023). Yet integrating personalization into the end-to-end paradigm introduces unique challenges. Unlike modular systems operating within narrowly defined scenarios, E2EAD requires preference generalization across a broad spectrum of complex, real-world scenarios. Meeting this demand necessitates large-scale datasets that not only cover diverse traffic conditions but also provide well-crafted driving style annotations.
@@
nuScenes Caesar et al. (2020) Real City OL
OpenScene Peng et al. (2023) Real City&Rural OL
Longest6 Chitta et al. (2022) Sim City CL
CARLA Contributors (2024) Sim City CL
MetaDrive Li et al. (2022) Sim City CL
Bench2Drive Jia et al. (2024) Sim City CL
NAVSIM Dauner et al. (2024) Real City&Rural Semi-CL
HITL-RampMerging Li et al. (2023) HITL Ramp-Merge OL
HITL-CarFollowing Zhao et al. (2022) HITL Car-Following OL
HITL-LaneChange Liao et al. (2023) HITL Lane-Change OL
HITL-MultiScene Ke et al. (2024) HITL City OL
UAH Romera et al. (2016) Real City&highway OL
Brain4Cars Jain et al. (2016) Real Lane-Change&Merge OL
PDB Wei et al. (2025) Real City OL
@@ 2 Related Work
Personalization has long been recognized as a key factor to enhance user comfort, trust, and acceptance of AD Liao et al. (2024). In conventional modular pipelines, personalized strategies have been extensively explored, where the system can adapt to users’ preferences in stand-alone driving scenarios, such as car following, ramp merge, and lane-change behaviors Zhao et al. (2022); Li et al. (2023); Liao et al. (2023). Other early attempts personalize models in unrealistic human-in-the-loop (HITL) simulation Ke et al. (2024). However, these approaches often suffer from poor generalization in dynamic, real-world environments.
Recent advances in large language models (LLMs) have introduced new possibilities for personalization in AD, enabling the encoding of user intent and preferences to influence downstream decision-making Cui et al. (2024); Xu et al. (2024); Kou et al. (2025). Despite their potential, these methods remain constrained by modular frameworks and often depend on handcrafted features. MAVERIC Schrum et al. (2024) represents a notable early attempt to enable end-to-end personalized driving. Nevertheless, in the absence of standardized datasets and benchmarks, these methods face challenges in reproducibility and scalability. Despite these advancements, existing work seldom addresses how to merge user preferences in an E2EAD manner. A critical bottleneck is the lack of large-scale, real-world datasets with well-crafted annotations of personalized driving styles.
@@
End-to-end autonomous driving represents a promising paradigm that maps raw sensor inputs directly to driving actions or planned trajectories. Early work Codevilla et al. (2018) demonstrated the feasibility of E2EAD. Over the past decade, research has advanced toward more sophisticated architectures incorporating temporal reasoning, multimodal fusion, and Transformer-based methods Prakash et al. (2021); Jia et al. (2023); Shao et al. (2023).
Despite this progress, most E2EAD models are optimized for average-level objectives, lacking the capacity to adapt to user-specific preferences. While recent architectures improve generalization and interpretability Li et al. (2024); Zheng et al. (2024); Jia et al. (2025), they remain limited in capturing individualized driving styles. Moreover, the absence of dedicated datasets and standardized benchmarks for evaluating personalization continues to hinder systematic advancement in this area. To address these limitations, we propose a framework for learning style-conditioned E2EAD models and establish a benchmark to evaluate their alignment with human driving behavior.
@@
Benchmarking is fundamental to the development and evaluation of AD models. Typical AD datasets and benchmarks are summarized in Tab. 1. Early E2EAD benchmarks Caesar et al. (2020); Peng et al. (2023) adopt non-interactive open-loop settings with real-world data, but fail to capture behavioral feedback. Recent works have introduced closed-loop benchmarks in simulation, such as Longest6 Chitta et al. (2022), MetaDrive Li et al. (2022), CARLA Series Contributors (2024), which allow online policy evaluation but rely entirely on simulated scenarios. Semi-simulated platforms like NAVSIM Dauner et al. (2024) strike a balance by simulating behavior feedback based on real-world scenes. However, existing benchmarks remain focused on task performance and ignore user preferences, cannot evaluate personalized driving behavior.
Efforts to benchmark personalized AD have led to two main dataset categories: human-in-the-loop (HITL) simulations and real-world driving logs. 1) HITL datasets are typically collected in controlled scenarios such as ramp merging Li et al. (2023), car following Zhao et al. (2022), lane changing Liao et al. (2023), and multiple scenarios Ke et al. (2024). While these setups enable preference modeling, they often rely on simplified simulation engines and curated environments, introducing a domain gap between real-world conditions and unrealistic scenarios. 2) Real-world datasets such as UAH-DriveSet Romera et al. (2016), Brain4Cars Jain et al. (2016), and PDB Wei et al. (2025) provide behavioral annotations related to diving intent. These datasets have contributed substantially to understanding driver behavior and intent modeling. However, they are typically constrained to narrow driving contexts, such as highways or intersections. More importantly, they are not structured for end-to-end learning: most only provide high-level behavior labels without continuous perception-to-control supervision, and none include standardized evaluation protocols for comparing personalized driving policies.
@@ 3 StyleDrive Dataset
The StyleDrive dataset is sampled and constructed from the large-scale AD dataset OpenScene Peng et al. (2023), and comprises nearly 30k driving scenarios annotated with style preferences. OpenScene features trajectories collected from over 16 drivers across diverse urban and suburban environments in both Singapore and the USA. This diversity offers a rich foundation for labeling user-specific tendencies in various scenarios (also illustrated in Fig. 4). Beyond the original OpenScene data, we introduce a unified framework for modeling and annotating personalized driving preferences.
@@ 3.1 Framework for Modeling and Annotation of Driving Preference
To enrich the traffic scenario classification with dynamic, context-sensitive information, we apply a fine-tuned vision-language model (VLM) to extract high-level semantic cues from driving videos. Since generic VLMs lack task-specific perception and reasoning capabilities for AD, we adapt the Video-LLaMA3 model Zhang et al. (2025) via lightweight fine-tuning using the LingoQA dataset Marcu et al. (2024) - a multimodal dataset for reasoning about road semantics (more details in Sect.5 of the Supplementary).
@@ 4 StyleDrive Benchmark
Our benchmark is built on a lightweight simulator NavSim Dauner et al. (2024), which emphasizes non-reactive behavior simulation and grounded in real-world traffic scenarios. In our adaptation, the scenarios of StyleDrive are replayed, and style-conditioned E2EAD models learn policies in this playground.
@@ 4.1 Evaluation Metric: Style-Modulated PDMS
To evaluate whether a model ensures not only safety and feasibility but also alignment with a desired driving style, we propose the Style-Modulated Predictive Driver Model Score (SM-PDMS), an extension of the Predictive Driver Model Score (PDMS) introduced in the NavSim Dauner et al. (2024). SM-PDMS enhances the original PDMS by incorporating a behavioral alignment component that quantifies the degree to which a policy conforms to specified style preferences. Sub-metrics related to traffic safety and road adherence are retained without modification, as they are considered invariant across driving styles. In contrast, driving style is primarily manifested through motion dynamics, such as following distance, speed, and angular velocity, which in turn affect style-sensitive sub-metrics including comfort (Comf.), ego progress (EP), and time-to-collision (TTC).
@@ 4.2 Baseline Methods
  • AD-MLP-Style: Extends the AD-MLP baseline Zhai et al. (2023) with a one-hot driving style vector concatenated to ego states. The combined input is passed through MLPs for style-aware trajectory regression.
  • TransFuser-Style: Builds on TransFuser Chitta et al. (2022), which fuses image and LiDAR features for planning. A one-hot style vector is concatenated with the trajectory query, fused via an MLP to restore dimensionality, and then fed into the trajectory prediction head.
  • DiffusionDrive-Style: Modifies DiffusionDrive Liao et al. (2025) by injecting a one-hot style vector into the trajectory head. The vector is concatenated with agent features, fused via a fusion MLP, and followed by a regression MLP. This block runs twice for cascade refinement.
  • WoTE-Style: Adapts the BEV World model Li et al. (2025), which forecasts future BEV states. Driving style is injected into the offset prediction head via concatenation and MLP fusion, following the same strategy as in DiffusionDrive-Style.
@@
AD-MLP Zhai et al. (2023) 92.63 77.68 83.83 99.75 78.01 63.72
TransFuser Chitta et al. (2022) 96.74 88.43 91.08 99.65 84.39 78.12
WoTE Li et al. (2025) 97.29 92.39 92.53 99.13 76.31 79.56
DiffusionDrive Liao et al. (2025) 96.66 91.45 90.63 99.73 80.39 79.33
@@ 4.3 Main Results Analysis
WoTE Li et al. (2025) 0.733 1.434 2.349 1.506
AD-MLP Zhai et al. (2023) 0.503 1.262 2.383 1.382
TransFuser Chitta et al. (2022) 0.431 0.963 1.701 1.032
DiffusionDrive Liao et al. (2025) 0.471 1.086 1.945 1.167
@@ 1 Details about Style-aware Metric Design
The original PDMS framework Dauner et al. [2024] evaluates comfort by penalizing trajectories exhibiting excessive acceleration and jerk, relying on fixed thresholds derived from empirical studies or simulator dynamics. While this approach effectively discourages physically unstable motions, it applies a uniform standard to all driving styles, disregarding the fact that perceived comfort can vary substantially across driver types. For instance, what a conservative driver may consider abrupt, an aggressive driver might deem acceptable. By enforcing a static threshold regardless of context or preference, the metric overlooks stylistic nuance and risks penalizing behavior that is intentionally assertive yet still within acceptable human tolerances.
@@ 1.2 Ego Progress (EP)
In the original PDMS framework Dauner et al. [2024], the Ego Progress (EP) metric measures how far the ego vehicle advances along its predicted trajectory. This is done by comparing the evaluated agent’s cumulative progress to that of a fixed rule-based baseline agent (PDM) Dauner et al. [2023]. If the agent travels less distance than the PDM, the score is computed as the ratio of its progress to that of the PDM; otherwise, it is assigned a perfect score of 1. While this approach captures task completion in a basic sense, it introduces several limitations. First, it reduces progress evaluation to a binary assessment of whether an agent outperforms a single rule-based baseline. Second, it provides no notion of “how much” deviation is acceptable or stylistically consistent—agents that accelerate too aggressively or hesitate excessively are not penalized, as long as they surpass the PDM in distance. As a result, the original EP metric lacks sensitivity to human-like driving intent and cannot distinguish between stylistically misaligned trajectories.
@@ 1.4 Othe submetrics, Aggregation, and Notes
Other safety-related submetrics include NC and DAC defined in Navsim Dauner et al. [2024] are retained, and we also retain Navsim’s aggregation method for all submetrics.
The PDMS referenced throughout the paper and its appendix corresponds to the Navsim Dauner et al. [2024] version 1 implementation.
@@ 4.4 Heuristics in Crosswalk Scenarios
  • Conservative (C): If $v_{\text{avg}}$ is significantly low (e.g., <2.0<2.0< 2.0 m/s), regardless of other indicators.
@@ 5.1 Details of Benchmark Methods Training
Regarding optimization, the learning rate for both DiffusionDrive Liao et al. [2025] and DiffusionDrive-Style was set to 6e-4. For AD-MLPZhai et al. [2023], Transfuser Chitta et al. [2022], WoTE Li et al. [2025], as well as their style-conditioned variants (AD-MLP-Style, Transfuser-Style, WoTE-Style), the learning rate was consistently set to 1e-4.
@@ 5.2 Details of VLM Fine-tuning towards Better Driving Scene Understanding
To enable accurate and context-sensitive semantic reasoning in driving scenes, we fine-tune the Video-LLaMA3 model Zhang et al. [2025] using lightweight parameter-efficient tuning (LoRA). This process tailors the model to better understand autonomous driving scenarios, which are underrepresented in generic vision-language pretraining.
@@
We build upon the LingoQA dataset Marcu et al. [2024], a multimodal benchmark comprising video clips and scene-level question-answer pairs focused on road semantics, interactions, and intent recognition. The dataset includes approximately 267.8k video–prompt–response triplets, covering diverse topology and behavioral cases.

### v3

@@ 1 Introduction
As autonomous driving (AD) technologies continue to mature, one of the key factors for translating these systems into real-world products is personalization. Tailoring vehicle behavior to individual user preferences is essential for enhancing user experience, building trust, and fostering long-term adoption hasenjager2017personalization. Traditional modular systems have extensively explored personalized strategies, enabling adaptations to specific driving preferences. However, these methods often rely on isolated, scenario-specific adaptations tian2022personalized or unrealistic human-in-the-loop (HITL) simulation ke2024d2e, which severely limit their generalization to dynamic, real-world environments. Furthermore, the fragmented nature of modular systems zhu2018personalized; cui2024personalized; kou2025padriver hinders scalability to massive real-world data. Due to these limitations, personalization remains largely underexplored in end-to-end autonomous driving (E2EAD), where perception, planning, and control are integrated in an unified architecture codevilla2018end. This gap presents a significant barrier to realizing human-centric AD at scale.
@@
This gap is particularly pressing. As E2EAD systems become increasingly capable and are deployed across diverse, open-world driving environments, aligning vehicle behavior with user preferences becomes essential for enhancing comfort, perceived safety, and long-term user acceptance speidel2019towards; aledhari2023motion. Yet integrating personalization into the end-to-end paradigm introduces unique challenges. Unlike modular systems operating within narrowly defined scenarios, E2EAD requires preference generalization across a broad spectrum of complex, real-world scenarios. Meeting this demand necessitates large-scale datasets that not only cover diverse traffic conditions but also provide well-crafted driving style annotations.
@@
nuScenes caesar2020nuscenes Real City OL
OpenScene peng2023openscene Real City&Rural OL
Longest6 chitta2022transfuser Sim City CL
CARLA carla_leaderboard_2024 Sim City CL
MetaDrive li2022metadrive Sim City CL
Bench2Drive jia2024bench2drive Sim City CL
NAVSIM dauner2024navsim Real City&Rural Semi-CL
HITL-RampMerging li2023personalized HITL Ramp-Merge OL
HITL-CarFollowing zhao2022personalized HITL Car-Following OL
HITL-LaneChange liao2023driver HITL Lane-Change OL
HITL-MultiScene ke2024d2e HITL City OL
UAH romera2016need Real City&highway OL
Brain4Cars jain2016brain4cars Real Lane-Change&Merge OL
PDB wei2025pdb Real City OL
@@ 2 Related Work
Personalization has long been recognized as a key factor to enhance user comfort, trust, and acceptance of AD liao2024review. In conventional modular pipelines, personalized strategies have been extensively explored, where the system can adapt to users’ preferences in stand-alone driving scenarios, such as car following, ramp merge, and lane-change behaviors zhao2022personalized; li2023personalized; liao2023driver. Other early attempts personalize models in unrealistic human-in-the-loop (HITL) simulation ke2024d2e. However, these approaches often suffer from poor generalization in dynamic, real-world environments.
Recent advances in large language models (LLMs) have introduced new possibilities for personalization in AD, enabling the encoding of user intent and preferences to influence downstream decision-making cui2024personalized; xu2024drivegpt4; kou2025padriver. Despite their potential, these methods remain constrained by modular frameworks and often depend on handcrafted features. MAVERIC schrum2024maveric represents a notable early attempt to enable end-to-end personalized driving. Nevertheless, in the absence of standardized datasets and benchmarks, these methods face challenges in reproducibility and scalability. Despite these advancements, existing work seldom addresses how to merge user preferences in an E2EAD manner. A critical bottleneck is the lack of large-scale, real-world datasets with well-crafted annotations of personalized driving styles.
@@
End-to-end autonomous driving represents a promising paradigm that maps raw sensor inputs directly to driving actions or planned trajectories. Early work codevilla2018end demonstrated the feasibility of E2EAD. Over the past decade, research has advanced toward more sophisticated architectures incorporating temporal reasoning, multimodal fusion, and Transformer-based methods prakash2021multi; jia2023think; shao2023safety.
Despite this progress, most E2EAD models are optimized for average-level objectives, lacking the capacity to adapt to user-specific preferences. While recent architectures improve generalization and interpretability li2024enhancing; zheng2024genad; jia2025drivetransformer, they remain limited in capturing individualized driving styles. Moreover, the absence of dedicated datasets and standardized benchmarks for evaluating personalization continues to hinder systematic advancement in this area. To address these limitations, we propose a framework for learning style-conditioned E2EAD models and establish a benchmark to evaluate their alignment with human driving behavior.
@@
Benchmarking is fundamental to the development and evaluation of AD models. Typical AD datasets and benchmarks are summarized in Tab. 1. Early E2EAD benchmarks caesar2020nuscenes; peng2023openscene adopt non-interactive open-loop settings with real-world data, but fail to capture behavioral feedback. Recent works have introduced closed-loop benchmarks in simulation, such as Longest6 chitta2022transfuser, MetaDrive li2022metadrive, CARLA Series carla_leaderboard_2024, which allow online policy evaluation but rely entirely on simulated scenarios. Semi-simulated platforms like NAVSIM dauner2024navsim strike a balance by simulating behavior feedback based on real-world scenes. However, existing benchmarks remain focused on task performance and ignore user preferences, cannot evaluate personalized driving behavior.
Efforts to benchmark personalized AD have led to two main dataset categories: human-in-the-loop (HITL) simulations and real-world driving logs. 1) HITL datasets are typically collected in controlled scenarios such as ramp merging li2023personalized, car following zhao2022personalized, lane changing liao2023driver, and multiple scenarios ke2024d2e. While these setups enable preference modeling, they often rely on simplified simulation engines and curated environments, introducing a domain gap between real-world conditions and unrealistic scenarios. 2) Real-world datasets such as UAH-DriveSet romera2016need, Brain4Cars jain2016brain4cars, and PDB wei2025pdb provide behavioral annotations related to diving intent. These datasets have contributed substantially to understanding driver behavior and intent modeling. However, they are typically constrained to narrow driving contexts, such as highways or intersections. More importantly, they are not structured for end-to-end learning: most only provide high-level behavior labels without continuous perception-to-control supervision, and none include standardized evaluation protocols for comparing personalized driving policies.
@@ 3 StyleDrive Dataset
The StyleDrive dataset is sampled and constructed from the large-scale AD dataset OpenScene peng2023openscene, and comprises nearly 30k driving scenarios annotated with style preferences. OpenScene features trajectories collected from over 16 drivers across diverse urban and suburban environments in both Singapore and the USA. This diversity offers a rich foundation for labeling user-specific tendencies in various scenarios (also illustrated in Fig. 4). Beyond the original OpenScene data, we introduce a unified framework for modeling and annotating personalized driving preferences.
@@ 3.1 Framework for Modeling and Annotation of Driving Preference
To enrich the traffic scenario classification with dynamic, context-sensitive information, we apply a fine-tuned vision-language model (VLM) to extract high-level semantic cues from driving videos. Since generic VLMs lack task-specific perception and reasoning capabilities for AD, we adapt the Video-LLaMA3 model zhang2025videollama via lightweight fine-tuning using the LingoQA dataset marcu2024lingoqa - a multimodal dataset for reasoning about road semantics (more details in Sect.5 of the Supplementary).
@@ 4 StyleDrive Benchmark
Our benchmark is built on a lightweight simulator NavSim dauner2024navsim, which emphasizes non-reactive behavior simulation and grounded in real-world traffic scenarios. In our adaptation, the scenarios of StyleDrive are replayed, and style-conditioned E2EAD models learn policies in this playground.
@@ 4.1 Evaluation Metric: Style-Modulated PDMS
To evaluate whether a model ensures not only safety and feasibility but also alignment with a desired driving style, we propose the Style-Modulated Predictive Driver Model Score (SM-PDMS), an extension of the Predictive Driver Model Score (PDMS) introduced in the NavSim dauner2024navsim. SM-PDMS enhances the original PDMS by incorporating a behavioral alignment component that quantifies the degree to which a policy conforms to specified style preferences. Sub-metrics related to traffic safety and road adherence are retained without modification, as they are considered invariant across driving styles. In contrast, driving style is primarily manifested through motion dynamics, such as following distance, speed, and angular velocity, which in turn affect style-sensitive sub-metrics including comfort (Comf.), ego progress (EP), and time-to-collision (TTC).
@@ 4.2 Baseline Methods
  • AD-MLP-Style: Extends the AD-MLP baseline zhai2023rethinking with a one-hot driving style vector concatenated to ego states. The combined input is passed through MLPs for style-aware trajectory regression.
  • TransFuser-Style: Builds on TransFuser chitta2022transfuser, which fuses image and LiDAR features for planning. A one-hot style vector is concatenated with the trajectory query, fused via an MLP to restore dimensionality, and then fed into the trajectory prediction head.
  • DiffusionDrive-Style: Modifies DiffusionDrive liao2024diffusiondrive by injecting a one-hot style vector into the trajectory head. The vector is concatenated with agent features, fused via a fusion MLP, and followed by a regression MLP. This block runs twice for cascade refinement.
  • WoTE-Style: Adapts the BEV World model li2025end, which forecasts future BEV states. Driving style is injected into the offset prediction head via concatenation and MLP fusion, following the same strategy as in DiffusionDrive-Style.
@@
AD-MLP zhai2023rethinking 92.63 77.68 83.83 99.75 78.01 63.72
TransFuser chitta2022transfuser 96.74 88.43 91.08 99.65 84.39 78.12
WoTE li2025end 97.29 92.39 92.53 99.13 76.31 79.56
DiffusionDrive liao2024diffusiondrive 96.66 91.45 90.63 99.73 80.39 79.33
@@ 4.3 Main Results Analysis
WoTE li2025end 0.733 1.434 2.349 1.506
AD-MLP zhai2023rethinking 0.503 1.262 2.383 1.382
TransFuser chitta2022transfuser 0.431 0.963 1.701 1.032
DiffusionDrive liao2024diffusiondrive 0.471 1.086 1.945 1.167
@@ 1 Details about Style-aware Metric Design
The original PDMS framework dauner2024navsim evaluates comfort by penalizing trajectories exhibiting excessive acceleration and jerk, relying on fixed thresholds derived from empirical studies or simulator dynamics. While this approach effectively discourages physically unstable motions, it applies a uniform standard to all driving styles, disregarding the fact that perceived comfort can vary substantially across driver types. For instance, what a conservative driver may consider abrupt, an aggressive driver might deem acceptable. By enforcing a static threshold regardless of context or preference, the metric overlooks stylistic nuance and risks penalizing behavior that is intentionally assertive yet still within acceptable human tolerances.
@@ 1.2 Ego Progress (EP)
In the original PDMS framework dauner2024navsim, the Ego Progress (EP) metric measures how far the ego vehicle advances along its predicted trajectory. This is done by comparing the evaluated agent’s cumulative progress to that of a fixed rule-based baseline agent (PDM) dauner2023parting. If the agent travels less distance than the PDM, the score is computed as the ratio of its progress to that of the PDM; otherwise, it is assigned a perfect score of 1. While this approach captures task completion in a basic sense, it introduces several limitations. First, it reduces progress evaluation to a binary assessment of whether an agent outperforms a single rule-based baseline. Second, it provides no notion of “how much” deviation is acceptable or stylistically consistent—agents that accelerate too aggressively or hesitate excessively are not penalized, as long as they surpass the PDM in distance. As a result, the original EP metric lacks sensitivity to human-like driving intent and cannot distinguish between stylistically misaligned trajectories.
@@ 1.4 Othe submetrics, Aggregation, and Notes
Other safety-related submetrics include NC and DAC defined in Navsim dauner2024navsim are retained, and we also retain Navsim’s aggregation method for all submetrics.
The PDMS referenced throughout the paper and its appendix corresponds to the Navsim dauner2024navsim version 1 implementation.
@@ 4.4 Heuristics in Crosswalk Scenarios
  • Conservative (C): If $v_{\text{avg}}$ is significantly low (e.g., <2.0<2.0 m/s), regardless of other indicators.
@@ 5.1 Details of Benchmark Methods Training
Regarding optimization, the learning rate for both DiffusionDrive liao2024diffusiondrive and DiffusionDrive-Style was set to 6e-4. For AD-MLPzhai2023rethinking, Transfuser chitta2022transfuser, WoTE li2025end, as well as their style-conditioned variants (AD-MLP-Style, Transfuser-Style, WoTE-Style), the learning rate was consistently set to 1e-4.
@@ 5.2 Details of VLM Fine-tuning towards Better Driving Scene Understanding
To enable accurate and context-sensitive semantic reasoning in driving scenes, we fine-tune the Video-LLaMA3 model zhang2025videollama using lightweight parameter-efficient tuning (LoRA). This process tailors the model to better understand autonomous driving scenarios, which are underrepresented in generic vision-language pretraining.
@@
We build upon the LingoQA dataset marcu2024lingoqa, a multimodal benchmark comprising video clips and scene-level question-answer pairs focused on road semantics, interactions, and intent recognition. The dataset includes approximately 267.8k video–prompt–response triplets, covering diverse topology and behavioral cases.